OpenStudy (anonymous):

A system of equations is shown below: 6x = -7y + 11 5x + 2y = 13 What is the solution to this system of equations? (-3, 1) (-1, 3) (1, -3) (3, -1)

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Solve the following system: y = x + 3 2x + y = 9 (2, 5) (5, 2) (-2, 5) (2, -5)

OpenStudy (anonymous):

The graph below shows the solution to a system of inequalities: Solid line joining ordered pairs 0, 3.75 and 15, 0. Shade the portion of the graph above the line in the first and second quadrant Which of the following inequalities is modeled by the graph? 5x + 4y ≥ 15; y ≥ 0 5x - 4y ≥ 15; y ≥ 0 5x + 4y ≤ 15; y ≥ 0 -5x - 4y ≥ 15; y ≥ 0

DivineSolar (divinesolar):

@alyssa_fields50 second one is C